# Generalizing Deepfake Video Detection with Plug-and-Play: Video-Level Blending and Spatiotemporal Adapter Tuning

### v1

@@ 4.6 Ablation Study
Can image blending method, e.g., SBI, act as a “video augmentor"? We designed an ablation study to investigate the effectiveness of fixed SBI (without temporal inconsistency) and VB (without spatial artifacts) in video detection. Results in Table 6 show that SBI achieves lower results than VB, but the combination of both SBI and VB is more effective than VB alone. This suggests that SBI and VB could be complementary, with one focusing on spatial aspects and the other on temporal aspects. This finding may inspire future research on how to effectively utilize both techniques to improve generalization in video detection.
@@ 5 Conclusion
In this paper, we uncover a novel and previously underexplored temporal forgery artifact present across various manipulated videos, which we term Facial Feature Drift (FFD). To simulate this artifact, we propose a novel Video-level Blending data (VB), designed to aid the model in learning more general forgery artifacts within the temporal domain. We also propose an adapter-based strategy for efficiently and effectively learning both the spatial and temporal forgery artifacts, without being overfitting to one specific type (e.g., learning spatial only). Extensive experiments on several widely used and latest deepfake datasets verify the effectiveness of the proposed methods. We hope that our work will inspire further research into (1) the development of video augmentations (also the approach to combine both image and video blending), and (2) efficient spatiotemporal adapter tunning in the deepfake video detection community.

### v2

@@ 4.2 Generalization Performance Evaluation
4.3 Ablation Study
Effectiveness of our plug-and-play strategy.
Effectiveness of VB and StA.
@@ 4.3 Ablation Study
Table 9: Ablation studies of the proposed cross-attention.
Fusion Method CDF-v2 DFDC WDF DFo Avg.
Linear combination ($\mathbf{e}_{k}^{s}+\mathbf{e}_{k}^{t})$ 0.886 0.787 0.805 0.974 0.863
Cross-Attn ($\mathbf{S2T}^{\text{orig}}+\mathbf{T2S}^{\text{orig}}$) 0.947 0.843 0.848 0.991 0.905
@@
Can image blending method, e.g., SBI, act as a “video augmentor”? We designed an ablation study to investigate the effectiveness of fixed SBI (without temporal inconsistency) and VB (without spatial artifacts) in video detection. Results in Table 8 show that SBI achieves lower results than VB, but the combination of both SBI and VB is more effective than VB alone. This suggests that SBI and VB could be complementary, with one focusing on spatial aspects and the other on temporal aspects. This finding may inspire future research on how to effectively utilize both techniques to improve generalization in video detection.
Effectiveness of the proposed cross-attention module.
By our experiments, we surprisingly found that the cross-attention module is significantly important in our StA architecture (see Tab. 9). This might indicate that the spatial and temporal information can be benefited to each other, potentially in a high-level manner (not linear).
Evidence for the proposed spatiotemporal adapters can indeed learn spatial and temporal clues?
To verify it with visualizations, we create a (boring) video by repeating a single image into a video sequence. We then pass the input to our spatial (S) and temporal (T) adapters (two pathways with different kernel sizes of 3D Convs). As expected, Fig. 4 shows that the outputs of S remain constant (same inputs), while the outputs of T are zero (no change in time). Plus, for a normal video as input, T_adapter can capture reasonable temporal-related motion, e.g., mouth movement.
Comparison of different video-level blending solutions.
To reproduce and simulate the FFD artifact, we propose a novel video-level blending data called VB. We also attempt to use other choices for implementation. In this subsection, we discuss the effectiveness of different video-level blending solutions for reproducing the FFD. Here, we discuss three potential solutions: (1) CBI (blend with other frames within one clip); (2) PFIG, the same method proposed in [39] that performs the facial-region blending of two frames from different videos. First, we consider CBI, which performs the blending operation with another frame within the same clip. This method could introduce the inconsistent time shift of the facial region between frames. However, as we can see in Fig. 5, the inconsistency between frames is also too obvious, which does not align with the original fake (FFD). Second, we also consider PFIG [39], which performs blending between two frames (one real and one fake). PFIG randomly chooses one facial region of the fake and blends it into the real. Here, to simulate the FFD, we selectively choose a facial region from another frame within the same clip and blend it back into the real to generate the synthetic data. As we can see in Fig. 5, the blending region is mainly within a specific region (e.g., eyes), not aligned with the original fake.
@@ 5 Conclusion
In this paper, we uncover a novel and previously underexplored temporal forgery artifact present across various manipulated videos, which we term Facial Feature Drift (FFD). To simulate this artifact, we propose a novel Video-level Blending data (VB), designed to aid the model in learning more general forgery artifacts within the temporal domain. We also propose an adapter-based strategy for efficiently and effectively learning both the spatial and temporal forgery artifacts, without being overfitting to one specific type (e.g., learning spatial only). Extensive experiments on several widely used and latest deepfake datasets verify the effectiveness of the proposed methods.